### v1

# Trilinos: Enabling Scientific Computing Across Diverse Hardware Architectures at Scale

Abstract.
Trilinos is a community-developed, open-source software framework that facilitates building large-scale, complex, multiscale, multiphysics simulation code bases for scientific and engineering problems. Since the Trilinos framework has undergone substantial changes to support new applications and new hardware architectures, this document is an update to “An Overview of the Trilinos project” by Heroux et al. (ACM Transactions on Mathematical Software, 31(3):397–423, 2005). It describes the design of Trilinos, introduces its new organization in product areas, and highlights established and new features available in Trilinos. Particular focus is put on the modernized software stack based on the Kokkos ecosystem to deliver performance portability across heterogeneous hardware architectures. This paper also outlines the organization of the Trilinos community and the contribution model to help onboard interested users and contributors.
Heterogeneous Hardware Architectures; High-Performance Computing; Performance Portability; Scientific Software Frameworks
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Mathematics of computing Solversccs: Mathematics of computing Mathematical software performanceccs: Computing methodologies Massively parallel algorithmsccs: Software and its engineering Software libraries and repositoriesccs: Software and its engineering Software performance
1. Introduction
Trilinos is a community-driven open-source C++ software framework and collection of reusable scientific libraries designed to enable the development of scalable, high-performance algorithms for solving complex, multiscale, and multiphysics engineering and scientific problems on advanced computing architectures. While Trilinos can run on a variety of hardware platforms ranging from small workstations to large supercomputers, the typical use of Trilinos is on leadership-class systems with new or emerging hardware architectures.
Trilinos was originally conceived as a framework of three packages for distributed memory systems. The original Trilinos publication (Heroux et al., 2005) described the motivation, philosophy, and capabilities of Trilinos at that time. Seven years later, a two-part collection of 16 papers provided a new overview of Trilinos and many of its packages (Heroux and Willenbring, 2012; Heroux and Rouson, 2012a, b; Pawlowski et al., 2012a; Oldfield et al., 2012; Bochev et al., 2012; Boman et al., 2012; Baker and Heroux, 2012; Edwards et al., 2012; Pawlowski et al., 2012b; Spotz, 2012; Long et al., 2012; Morris et al., 2012; Howle et al., 2012; Bavier et al., 2012; Gaidamour et al., 2012) and described the intervening expansion of Trilinos’ capabilities and strategic goals for Trilinos. Trilinos today hews to the overall vision from two decades ago, but has evolved alongside changes in programming models, application needs, hardware architectures, and algorithms. Trilinos has grown from a library of three packages to more than fifty packages with functionality and features supporting a wide range of applications.
For much of its history, packages have relied heavily on the linear algebra classes provided by Epetra. Epetra does not support larger problems (2B+ unknowns), nor does it allow for accelerator offload. Since the packages of the old Epetra stack are scheduled to be archived and removed from the main Trilinos repository on GitHub by the end of 2025, this paper describes only the modern Trilinos software stack, which builds upon Tpetra and the Kokkos ecosystem to achieve performance portability across hardware architectures.
This article attempts to capture a snapshot the current state of Trilinos, as an update to the articles published thirteen and twenty years ago. Therefore it focuses on the major developments within Trilinos in the last decade as well as new features and functionality that have been added to advance scientific and engineering applications. It will give only an overview of the features, and we refer to the extensive list of references for the details of these features. We are also cognizant of the fact that this article describes software that is being actively developed and constantly evolving. Hence, we will focus on the high-level features and concepts that we expect to remain stable for several years. We also briefly touch upon the Trilinos community and software engineering issues with respect to Trilinos.
1.1. Modern Trilinos: Performance Portability through Kokkos
[…]
To avoid the necessity of diverging code paths, Trilinos leverages the Kokkos ecosystem [^2] https://kokkos.org (Trott et al., 2021) to write performance portable code. The Kokkos ecosystem originated from the native Trilinos packages Kokkos (Heroux, 2011) and Kokkos Kernels but split off a decade ago into a standalone project hosted and developed independently of Trilinos [^3] https://github.com/kokkos. Now, Trilinos provides snapshots of the two primary Kokkos subprojects (Core and Kernels) that reflect the latest release of Kokkos. Trilinos can also be built against version-compatible external installations of Kokkos — a capability required for interoperability with other Kokkos-based libraries.
Kokkos enables Trilinos developers to write single source implementations of their packages that perform well on all major HPC hardware architectures. Some of its design principles are also reflected in Trilinos API designs. In particular, the use of execution and memory spaces as parameters of data structures and algorithms is a common thread throughout Trilinos. This design characteristic enables Trilinos users to manage the heterogeneity of modern architectures and its impact on Trilinos data structures.
1.2. Trilinos Functionality
The features and capabilities of Trilinos are divided into several software units called packages. Each Trilinos package has a well-defined set of unique capabilities that is important for scientific or engineering applications. Packages are semi-autonomous, often having their own development team, users, and set of requirements and development principles. However, packages also follow a general set of Trilinos expectations such as having a designated point-of-contact and following software quality expectations (e.g., sufficient documentation, continuous integration testing, clearly defined dependencies, and using the Trilinos infrastructure for building and installation).
While most packages are developed natively in the Trilinos repository, some packages are developed externally and snapshotted into Trilinos. These are (in alphabetical order): Compadre [^4] https://github.com/sandialabs/compadre, Kokkos and Kokkos Kernels, Krino, SEACAS [^5] https://github.com/sandialabs/seacas, Sierra ToolKit (STK), and TriBITS [^6] https://github.com/TriBITSPub/TriBITS. Some reasons for snapshotting external packages into Trilinos include: external packages and repositories facilitate easy and independent usage by other software projects without having to pull in Trilinos; snapshotting provides an integration process for external packages into Trilinos, ensuring continual compatibility between all packages; and snapshotting allows easy integration into Trilinos-based application codes, alleviating users from managing these dependencies themselves. We refer to the Trilinos website (https://trilinos.github.io) for a brief description of these packages.
In the following sections of this paper, we group the Trilinos packages that share common objectives (e.g., solving linear systems) together into Trilinos’ five product areas: Core, Linear Solvers and Preconditioners, Nonlinear Solvers and Analysis Tools, Discretization Tools, and Framework Infrastructure. Trilinos’ packages and their assignment to a product area are illustrated in Figure 1.
Refer to captionCoreZoltan2ThyraTpetraTeuchosKokkos KernelsKokkosNonlinear Solversand Analysis ToolsPiroStokhosSacadoROLNOX & LOCADiscretization ToolsCompadrePanzerTempusPhalanxIntrepid2Linear Solvers and PreconditionersStratimikosAnasaziTekoShyLU & Amesos2MueLuFROSchIfpack2BelosFrameworkPyTrilinos2DocumentationBuilding & Testing
Figure 1. Organization of the Trilinos library into five product areas and their packages: Product areas are labeled in bold font with their respective packages. Snapshotted packages described in this paper are marked by dashed boxes.
We briefly describe these product areas below.
Core
Core packages cover all aspects of creating, distributing, and mapping data to processing elements (cores, threads, nodes), load balancing, and redistributing data. This includes Trilinos’ abstractions for linear algebra data structures and algorithms as well as concrete implementations such as Tpetra linear algebra data structures. On a modern accelerator-based compute node, the abstractions provided by the Kokkos library become critical for Tpetra. These capabilities are described in detail in Section 2.
Linear Solvers and Preconditioners
The wide variety of applications that use Trilinos need a diverse set of linear solvers. Trilinos has support for both iterative and direct linear solvers, including interfaces to external solver packages. There are a number of preconditioner options from multithreaded or performance portable node-level preconditioners to scalable multilevel domain decomposition or multigrid preconditioners. The preconditioners and solvers use the data abstractions from the core packages. Section 3 provides a detailed description of these features.
Nonlinear Solvers and Analysis Tools
These packages provide high-level algorithms for computational simulations and design. Capabilities include solvers for nonlinear equations, parameter continuation, bifurcation tracking, optimization, and uncertainty quantification. Trilinos also provides lower-level utility packages to evaluate quantities of interest required by the analysis algorithms. Capabilities include automatic differentiation technology to evaluate derivatives and embedded ensemble propagation for uncertainty quantification. These packages and their capabilities will be discussed further in Section 4.
Discretization Tools
This collection of packages provides functionality for the discretization of differential equations. In particular, it supports mesh-free and mesh-based spatial discretizations, with a focus on high-order finite elements, and time integration. Discretization tools also include cross-cutting utilities for algorithmic differentiation and for managing directed acyclic graphs of evaluation kernels. These capabilities are described further in Section 5 in detail.
Framework
The Framework product is different than the other Trilinos products in that most of the resources and services are not associated with Trilinos packages. The Framework product area is primarily focused on activities such as developing and maintaining infrastructure for automated testing and documentation, as well as associated workflows. A small number of infrastructure and cross-cutting packages are also associated with the Framework, for example the Python wrapper package PyTrilinos2. Section 6 provides further details.
1.3. Outline of this manuscript
Sections 26 provide details on Trilinos’ product areas. Section 7 briefly touches upon the Trilinos community, software engineering aspects as well as the integration of Trilinos into the landscape of scientific software frameworks. Finally, Section 8 concludes the manuscript.
2. Trilinos Core
The Core product area provides essential tools for managing and distributing data across processing elements in parallel computing environments, addressing challenges such as load balancing and data redistribution. Key components discussed here include the Kokkos performance portability framework, Teuchos tools, Tpetra distributed-memory linear algebra, Thyra abstractions for high-level formulations of numerical algorithms, and Zoltan2 for combinatorial algorithms in parallel computing.
2.1. Performance Portability: Kokkos Core
Kokkos Core [^7] For historical reasons the package name is “Kokkos” since it precedes the creation of other Kokkos subprojects such as Kokkos Kernels. is a programming model designed for performance portability across hardware architectures such as CPUs and GPUs. Two primary insights into achieving performance portability shaped Kokkos’ development: i) abstractions for both parallel execution and data structures are required, and ii) the fundamental programming model must be descriptive rather than prescriptive.
The Kokkos programming model includes execution spaces to specify where an algorithm should run, execution policies to indicate how the algorithm should be run in terms of hardware resources, and execution patterns to designate whether the for, reduce, or scan pattern will be used by the algorithm. The data structure abstractions are built around the concepts of memory spaces to indicate where memory is allocated, memory layouts to specify how data is laid out in memory, and memory traits that enable customizations of how data is accessed. The primary data structure incorporating these concepts is Kokkos::View, a multidimensional array abstraction that inspired the ISO C++23 mdspan class [^8] https://www.open-std.org/jtc1/sc22/wg21/docs/papers/2022/p0009r18.html. Detailed descriptions of Kokkos Core can be found in (Edwards et al., 2014) and (Trott et al., 2022).
2.2. Performance Portable Kernels: Kokkos Kernels
[…]
The implementations of Kokkos Kernels algorithms leverage the hierarchical parallelism exposed by the Kokkos library (Kim et al., 2017). To ensure flexibility for the distributed libraries that might call its algorithms, Kokkos Kernels provides thread-safe implementations for most of its kernels, with increasing coverage for asynchronous implementations that allow an execution space instance to be passed when calling a kernel. Kokkos Kernels also serves as a major point of integration for vendor optimized libraries such as cuBLAS, cuSPARSE, rocBLAS, rocSPARSE, MKL, ARMpl and others.
The capabilities provided by Kokkos Kernels can be divided into four major categories: 1. BLAS algorithms, 2. sparse linear algebra and preconditioners, 3. graph algorithms, and 4. batched dense and sparse linear algebra (Liegeois et al., 2023). The main points of integration of Kokkos Kernels in Trilinos are Tpetra for the dense and sparse linear algebra capabilities, Ifpack2 for the preconditioners and batched algorithms, and the multigrid package MueLu which relies directly and indirectly on these general capabilities as well as more specialized algorithms such as graph coloring/coarsening (Kelley and Rajamanickam, 2022) and fused Jacobi-SpGEMM (generalized sparse matrix-matrix multiply) kernels.
Similar to the Kokkos library, Kokkos Kernels is developed in its own GitHub repository [^9] https://github.com/kokkos/kokkos-kernels outside of the Trilinos GitHub repository. Every version of the library is integrated and tested in Trilinos as part of the Kokkos ecosystem release process. Additional information on Kokkos Kernels’ capabilities can be found in (Deveci et al., 2018; Wolf et al., 2017).
2.3. Tools: Teuchos
Teuchos provides a suite of common tools for many Trilinos packages. These tools include memory management classes (Bartlett, 2010) such as smart pointers and arrays, parameter lists for communicating hierarchical lists of parameters between library or application layers, templated wrappers for BLAS and LAPACK, XML parsers, MPI, and other utilities. They provide a unified “look and feel” across Trilinos packages and help avoid common programming mistakes.
2.4. Distributed-Memory Linear Algebra: Tpetra
Tpetra (Baker and Heroux, 2012; Hoemmen, 2015) provides the distributed-memory infrastructure for linear algebra objects such as sparse graphs, sparse matrices, and dense vectors. These linear algebra objects are implemented using MPI-based abstractions for the inter-process communication, and Kokkos-based data structures for the portions of the object owned by each process (MPI rank). Kokkos Kernels is used for performing local, parallel computations within each MPI rank.
The objects provided are templated on scalar type (e.g., double, float, or a Sacado or Stokhos type), local index type, global index type, and Kokkos device (pair of execution and memory spaces). The distribution of the linear algebra objects is described by a Map data structure, which defines which elements are present on which ranks. Here “element” could be a row or column of a sparse matrix or an entry of a vector. Global indices number elements across an entire distributed object, while local indices number elements only within a rank. Since memory capacity limits the sizes of local objects, local indices can be safely represented with a more compact type than global indices (e.g., 32-bit int instead of 64-bit long).
DistObject is Tpetra’s abstract base class for general distributed objects. Subclasses define how to pack elements into a buffer for sending with MPI, and how to unpack received elements. The Import and Export classes describe how one DistObject communicates data to another. They are both created from a pair of Maps, a source and a target. For example, halo exchange on a vector is an Import where the source is the vector’s original distribution, and the target is the same except that each rank also “owns” the degrees-of-freedom in the halo around its subdomain.
The following are the primary linear algebra objects in Tpetra (all are subclasses of DistObject):
  • (Multi)Vectors: storage of dense vectors or collections of vectors (multivectors) and associated BLAS-1 like kernels (e.g., dot products, norms, scaling, vector addition, pointwise vector multiplication) as well as tall skinny QR (TSQR) factorization for multivectors.
  • CrsGraph a sparse graph in compressed row storage (CRS) format. Graphs also include import/export objects for use in halo/boundary exchanges associated with the graph.
  • CrsMatrix and BlockCrsMatrix are sparse matrices in CRS and block compressed sparse row (BSR) formats. Associated kernels include sparse matrix-vector product (SPMV), sparse matrix-matrix multiplication (SPGEMM) and triple-product, sparse matrix-matrix addition, sparse matrix transpose, diagonal extraction, Frobenius norm calculation, and row/column scaling.
2.5. Abstract Numerical Algorithms: Thyra
The Thyra package provides abstractions for implementing high-level abstract numerical algorithms (ANAs). Thyra is used by other Trilinos packages for implementing ANAs such as solvers for linear and nonlinear equations, bifurcation and stability analysis, optimization, and uncertainty quantification. The abstractions hide concrete linear algebra and solver implementations so that the algorithms can be written once and reused across many application codes. For example, one could write a Krylov solver and use it with multiple concrete linear algebra implementations such as Tpetra in Trilinos or the vector/matrix objects in PETSc (Balay et al., 2021).
Thyra is organized into three sets of abstractions depending on the type of operation required. All are templated on the scalar type. These sets are:
  • Vector/Operator Interfaces: The first level is an abstraction for vectors and operators (Bartlett, 2007). This set includes abstractions for vector spaces, vectors, multivectors, and linear operators. Thyra provides a default set of vector-vector and vector-operator operations for standard linear algebra tasks such as AXPY from BLAS. Defining an extensible interface such that users can add arbitrary vector/matrix operations and maintain performance can be difficult. A general tool to achieve this is the Trilinos utility package RTOp (Bartlett et al., 2004), which provides an interface for users to add new Thyra operations. RTOp accepts a vector of Thyra input objects and a vector of Thyra output objects along with a functor. RTOp will then apply the functor to each element of the data.
  • Operator Solve Interfaces: The second set of abstractions supplies abstract linear solvers that can be attached to operators. The interfaces include base classes for preconditioners, preconditioner factories, linear operators with a solver, and factories for linear operators with a solver. The factories build and update the preconditioners and linear solvers as needed. The solver interfaces allow one to wrap any concrete linear solver or preconditioner for use with the ANAs. Trilinos provides a separate library named Stratimikos that wraps all of the Trilinos preconditioners and linear solvers in Thyra abstractions. Run-time configuration of Stratimikos is driven by Teuchos::ParameterList options. Details on Stratimikos are given in Section 3.8.
  • Nonlinear Interfaces: The final set of abstractions supports nonlinear operations. Thyra provides an interface that wraps nonlinear solvers. Implicit time integration and optimization solvers typically require nonlinear solves and can abstract away the concrete implementation via Thyra. The NOX package in Trilinos (described in Section 4.1) provides a concrete implementation of this interface. The final interface in Thyra is for querying a physics application for common objects required by high-level Newton-based solvers. The Thyra::ModelEvaluator interface allows codes to ask the application to evaluate residuals, Jacobians, parametric sensitivities, Hessians, and post-processed quantities/derivatives of interest. The interface is designed to be stateless (i.e., consecutive calls with the same inputs should produce the same outputs). This design was chosen to avoid pitfalls in complex application codes interacting with general nonlinear algorithms. The design does not preclude its use in stateful applications that have a “memory” between timesteps, but rather forces the explicit use of the stateful information in the interface.
2.6. Combinatorial algorithms with hybrid parallelism: Zoltan2
Zoltan2 is a package for partitioning, load balancing, and other combinatorial scientific computing. It can be viewed as a successor to Zoltan (Boman et al., 2012) that provides both tighter integration with other Trilinos packages and support for hybrid parallelism. It is used primarily to partition and distribute data objects (such as sparse matrices) across processors.
Zoltan2 has a highly modular design. The user provides input via input adapters. Adapters for sparse graphs and matrices are provided for Xpetra, which can also be used to interface Tpetra data structures. The models create a graph, hypergraph, or coordinate model from the user input, which is typically a sparse matrix, mesh, or vector. The algorithms then use this model to solve the particular problem (e.g., partitioning or coloring) requested by the user.
Zoltan2 supports both distributed-memory parallelism and on-node parallelism for both CPU and GPU (via Kokkos). Algorithms that support such hybrid parallelism include Multi-jagged (Deveci et al., 2016) (geometric partitioning), Sphynx (Acer et al., 2021) (spectral graph partitioning), graph coloring (Bogle et al., 2020), and MPI task placement (Deveci et al., 2019).
3. Linear Solvers and Preconditioners
Trilinos offers many linear solver capabilities: dense and sparse direct solvers, iterative solvers, shared-memory preconditioners local to a compute node, and scalable distributed memory domain decomposition and multigrid methods. Furthermore, interfaces to several third-party direct solvers are provided. All native solver capabilities are built on top of Kokkos and are GPU capable to varying degrees; any exceptions are noted in the detailed descriptions below.
3.1. Iterative Linear Solvers and Krylov Methods: Belos
Belos (Bavier et al., 2012) provides next-generation iterative linear solvers and a powerful developer framework for solving linear systems and least-squares problems. This framework provides several abstractions that facilitate code reuse and extensibility.
First, Belos decouples the algorithms from the implementation of the underlying linear algebra object using traits mechanisms. While concrete linear algebra adapters are provided for Tpetra and Thyra, users can also implement their own interfaces to leverage any existing matrix and vector representations. Second, there are abstractions to orthogonalization to ease the integration of application or architecture-specific orthogonalization methods. Implementations of iterated classical Gram-Schmidt, DGKS-corrected (Daniel, Gragg, Kaufman, and Stewart-corrected) classical Gram-Schmidt, and iterated modified Gram-Schmidt algorithms are provided. Finally, powerful solver managers encapsulate the strategy for solving a linear system or least-squares problem using abstract interfaces to iteration kernels. As a result, the algorithms developed using Belos abstractions can be relatively agnostic to data layout in memory or distributed over processors and parallel matrix/vector operations.
Belos supersedes the AztecOO package (Heroux, 2004) providing solver managers for single-vector iterative solvers, but also extensions of these methods to block iterative methods for solving linear systems with multiple right-hand sides. Single-vector solver managers include conjugate gradient (CG), minimum residual (MINRES), generalized minimal residual (GMRES), stabilized biconjugate gradient (BiCGStab), and transpose-free QMR (TFQMR) as well as “seed” solvers (hybrid GMRES, PCPG), subspace recycling solvers (GCRO-DR and RCG), and least-squares solvers (LSQR). There are versions of CG, GMRES, and GCRO-DR that construct a block Krylov subspace to solve for one or more right-hand sides. There are also “pseudo-block” versions of CG, GMRES, and TFQMR that apply the single-vector iteration simultaneously on multiple right-hand sides, where matrix and preconditioner applications are aggregated to achieve better performance. Any single-vector iteration can be used to solve multiple right-hand sides as well but, if a block or pseudo-block version is not used, the solver will sequentially solve the right-hand sides.
In recent years, application or architecture-specific variants of the classic iterative methods have been integrated into Belos. Some of these variations are so minor that they are enabled through input parameters on the classic method. For instance, flexible GMRES (Saad, 1993) is an option for the block GMRES solver, and pipelined CG (Ghysels and Vanroose, 2014) is an option for both the block CG and pseudo-block CG solver. Additionally, there are stand-alone solvers for a pseudo-block stochastic CG method (Parker and Fox, 2012) and a fixed-point iteration method.
3.2. One-Level Domain Decomposition and Basic Iterative Methods: Ifpack2
Trilinos provides domain decomposition approaches in two different packages: Ifpack2 and ShyLU (specifically the ShyLU_DD and FROSch subpackages). Ifpack2 implements overlapping additive Schwarz approaches with several options for the local subdomain solves. The local subdomain solvers may be either CPU-only versions of incomplete factorization preconditioners implemented in Ifpack2 itself, such as ILU(k) and ILUt (thresholded ILU), or architecture-portable algorithms for incomplete factorizations and triangular solvers implemented in Kokkos Kernels. It is also possible to use direct solvers as subdomain solvers, and ShyLU also supports shared-memory inexact incomplete factorization preconditioners. Such one-level preconditioners can be used as smoothers within multigrid methods, for solving “simpler” problems where the setup cost of more robust and scalable multilevel methods is prohibitive in comparison to the reduction in the number of iterations, or when the underlying problem is simply not amenable to multilevel methods.
[…]
3.3. Multilevel Domain Decomposition Methods: FROSch
FROSch (Fast and Robust Overlapping Schwarz) is a framework for the construction of multilevel Schwarz domain decomposition preconditioners. Besides parallel scalability, FROSch emphasizes applicability and robustness across a wide range of challenging problems while supporting an algebraic construction. Specifically, most preconditioners can be built using only the fully assembled system matrix, though some variants can take advantage of additional geometric inputs. The algebraic construction relies on two key algorithmic components: first, the creation of an overlapping domain decomposition at the initial level based on the sparsity pattern of the system matrix, similar to Ifpack2; and second, the integration of extension-based coarse spaces, as in the classical two-level Generalized Dryja–Smith–Widlund (GDSW) preconditioner (Dohrmann et al., 2008) and its variants. While the initial version of FROSch (Heinlein et al., 2016) was based on the outdated Epetra linear algebra framework, the current implementation (Heinlein et al., 2020) leverages Xpetra. Originally designed as a lightweight wrapper around Epetra and Tpetra, the Xpetra package facilitated over the years compatibility with both the Epetra and Tpetra stacks. With the deprecation of Epetra, it now exclusively supports the Tpetra stack. Algorithmic variants of Schwarz methods implemented in FROSch include:
  • Extension-based coarse spaces based on a partition of unity on the interface, such as classical GDSW, reduced dimension GDSW (RGDSW) coarse spaces, and multiscale finite element method (MsFEM) coarse spaces; cf. (Heinlein et al., 2016, 2018);
  • Monolithic Schwarz preconditioners for block systems; cf. (Heinlein et al., 2019a).
  • Multilevel Schwarz preconditioners, which are obtained from two-level Schwarz preconditioners by recursively applying Schwarz preconditioners as an inexact solver for the coarse problems; cf. (Heinlein et al., 2022b).
FROSch has been applied to various challenging application problems, including: scalar elliptic and elasticity problems (Heinlein et al., 2016), possibly with heterogeneities (Alves et al., 2024); computational fluid dynamics problems (Heinlein et al., 2019a, [n. d.]); time-harmonic Maxwell’s and fluid-structure interaction problems (Heinlein et al., 2024); pharmaco-mechanical interactions in arterial walls (Balzani et al., [n. d.]); and coupled multiphysics problems for land ice simulations (Heinlein et al., 2022a). The latter three have been solved using monolithic preconditioning techniques. To extend robustness for heterogeneous model problems, an implementation of spectral coarse spaces (Heinlein et al., 2019b) is currently under development. FROSch preconditioners have scaled to more than $200,000$ cores on the Theta Cray XC40 supercomputer at the Argonne Leadership Computing Facility (ALCF); cf. (Heinlein et al., 2022b).
In its current implementation, FROSch assumes a one-to-one correspondence of subdomains and MPI ranks. Using an interface to the other solver packages in Trilinos, a variety of inexact solvers can be employed for the solution of the local subdomain problems. An extension to multiple subdomains per MPI rank is currently being implemented. Using Kokkos and Kokkos Kernels, FROSch has recently also been ported to GPUs (Yamazaki et al., 2023) with performance gains for the triangular solve or inexact solves with ILU on GPUs.
A demo/tutorial for FROSch can be found at the GitHub repository (fro, [n. d.]).
3.4. Multigrid Methods: MueLu
MueLu is a flexible and scalable high-performance multigrid solver library. It provides a variety of multigrid algorithms for problems ranging from Poisson-like operators, over elasticity, convection-diffusion, Navier-Stokes, and Maxwell’s equations, all the way to multigrid methods for coupled multiphysics systems. Besides its strong focus on aggregation-based algebraic multigrid (AMG) methods, MueLu comes with specialized capabilities for (semi-)structured grids to perform semi-coarsening along grid lines while forming the coarse operator via a Galerkin product (in contrast to classical geometric multigrid methods). MueLu is extensible and allows for the research and development of new multigrid preconditioning methods. Its weak and strong scalability up to $131,000$ cores of a Cray XC40 and one million cores of a Blue Gene/Q system, even for vector-valued partial differential equations (PDEs) on unstructured meshes, have been shown in (Lin et al., 2017; Thomas et al., 2019).
MueLu provides several approaches to constructing and solving the multilevel problem:
  • Algebraic smoothed aggregation approach (Vaněk et al., 1996): The matrix graph is colored to create aggregates (groups) of nodes. These aggregates define a tentative projection operator. A final projection operator is created by applying a smoother to the tentative operator. There are a variety of deterministic and non-deterministic coloring algorithms implemented directly in MueLu and in Kokkos Kernels. For a full description, see (Berger-Vergiat et al., 2023).
  • Algebraic multigrid for Maxwell’s equations: MueLu implements specialized solvers for the solution of curl-curl problems (Bochev et al., 2008). Scaling results on Haswell, KNL, ARM and NVIDIA V100 GPUs at full machine scale can be found in (Bettencourt et al., 2021).
[…]
  • Semi-coarsening algebraic multigrid approach (Tuminaro et al., 2016): Specialized aggregation procedures for three-dimensional meshes generated by extrusion of a two-dimensional unstructured grid allow to first coarsen in the direction of extrusion to reduce the system to a two-dimensional representation and then perform classical aggregation-based AMG for the remaining coarsenings.
  • AMG for (semi-)structured grids: MueLu has a structured aggregation capability in which the user may specify the coarsening rate used to compute interpolation operators. The coarse operators are then formed via a Galerkin product to avoid remeshing on the coarse levels. This work has been extended to semi-structured grids to leverage structured-grid computational performance also for globally unstructured grids (Mayr et al., 2022).
  • Geometric multigrid / matrix-free capabilities: MueLu treats the objects in its hierarchy as operators instead of matrices whenever possible. This enables one to use MueLu in a matrix-free fashion provided the user can supply their own grid transfers and coarse operators.
MueLu has a required dependency on Kokkos, and the main code paths in MueLu (e.g., smoothed aggregation, semi-coarsening, $p$-coarsening, structured aggregation) utilize Kokkos and have been shown to scale on device (Bettencourt et al., 2021). Some of the earlier algorithms such as energy minimization (Sala and Tuminaro, 2008) and the variable degree-of-freedom Laplacian have not been adapted to fully utilize Kokkos yet, however, they are part of ongoing efforts to use Kokkos throughout MueLu completely and to merge serial and device-capable algorithms.
There are a variety of resources to learn more about MueLu. An overview is given on the MueLu website [^10] https://trilinos.github.io/muelu.html. The MueLu User’s Guide (Berger-Vergiat et al., 2023) summarizes installation instructions and a reference to most of MueLu’s configuration parameters. The MueLu Tutorial (Mayr et al., 2023) introduces beginners and experts to various topics in MueLu and shows how to solve or precondition different linear systems using MueLu. Details on the compatibility of MueLu and its predecessor ML (Heroux et al., 2005; Gee et al., 2006) can be found in the MueLu User’s Guide (Berger-Vergiat et al., 2023).
Besides its C++ API, MueLu offers a MATLAB interface, MueMex, to provide access to MueLu’s aggregation and solver routines from MATLAB. MueMex allows users to setup and solve arbitrarily many problems with either MueLu as a preconditioner, Belos as a solver and Epetra or Tpetra for data structures.
MueLu requires the packages Teuchos, Tpetra, Xpetra, Kokkos, Kokkos Kernels, Amesos2, Ifpack2, and Zoltan2. Additionally, it supports interfaces to abstraction layer packages such as Stratimikos and Thyra through the MueLu adapters library. These interfaces are required to use MueLu with the Teko block preconditioning package. For more details on using MueLu with Teko, see the MueLu examples directory in Trilinos.
3.5. Direct Linear Solvers: ShyLU and Amesos2
ShyLU (Thornquist and Rajamanickam, 2014) provides an implementation of three solvers: Basker, Tacho, and a distributed-memory solver based on the Schur complement method.
Basker (Booth et al., 2017) is a sparse direct solver based on LU factorization for the problems that have the block triangular form (BTF) typically seen in circuit simulation applications. Basker uses this structure to factor and solve the diagonal blocks in parallel. The larger diagonal blocks can themselves be factored in parallel by discovering the parallelism available using a nested-dissection reordering. Basker focuses on exploiting thread-parallelism on the multi-core CPU architectures. Amesos2 also has a templated implementation of a sequential KLU solver called KLU2, which also exploits the BTF structure.
Tacho is a sparse direct solver that exploits supernodal block structure commonly found in sparse direct factorizations of matrices from mechanics applications. Tacho exploits this supernodal structure for both factorization and triangular-solve phases. It is based on Kokkos. Originally, Tacho implemented a task-parallel Cholesky factorization of sparse symmetric positive definite (SPD) matrices (Kim et al., 2018). However, to improve its portability, it has been extended to compute the sparse factorization based on level-set scheduling. Moreover, its functionality has been extended to the computation of an LDLT factorization of symmetric indefinite matrices, as well as the computation of an LU factorization of general matrices with a symmetric sparsity pattern.
In addition to their stand-alone use, the aforementioned node-level solvers may be used as the local solvers for domain decomposition preconditioners (Ifpack2 or FROSch) or as the coarse solvers for multilevel preconditioners (MueLu or FROSch).
ShyLU also implements a distributed-memory linear solver based on the Schur complement method (Thornquist and Rajamanickam, 2014). This hybrid approach combines direct and iterative solvers: each subdomain problem is solved in parallel using a direct solver, while the Schur complement is handled iteratively. The preconditioner for the Schur complement solver is computed using a probing approach or using a threshold-based dropping strategy. It has been developed to address the requirements of circuit simulation applications. The solver is also hybrid in terms of parallel computing, utilizing MPI+threads. However, it is not designed for GPU architectures, and therefore, it has not been implemented using Kokkos.
Amesos2 (Bavier et al., 2012) provides a uniform interface to direct linear solvers. These include third-party direct solvers such as CHOLMOD, MUMPS, Pardiso_MKL, SuperLU, SuperLU_MT, SuperLU_Dist, and STRUMPACK. Amesos2 also provides an interface to the on-node sparse direct solvers implemented in ShyLU.
3.6. Physics Block Operators and Preconditioners: Teko
The Teko library (Cyr et al., 2016) provides interfaces for operators and preconditioners that are constructed from large physics-based sub-blocks. The sub-blocks are Thyra operators which are themselves often implemented using Tpetra matrices. Teko provides generic block preconditioning strategies such as $2\times 2$ block-LU factorizations as well as block versions of Jacobi- and Gauss–Seidel-type methods for any number of blocks, as well as commonly used approximate inverse strategies for the Navier-Stokes equation such as SIMPLEC, LSC, and PCD preconditioners (Cyr et al., 2012). The Block-LU class offers the possibility to implement user-specific inverse approximation strategies, for example using a sparse approximate inverse to construct the Schur complement (Firmbach et al., 2024). More complicated multilevel hierarchies of block solvers can be generated via Teuchos::ParameterList objects. Block preconditioners for first-order formulations of Maxwell’s equations and Darcy flow are implemented in the Panzer package. Teko’s solvers can be registered in the Stratimikos interface for usage entirely driven by Teuchos::ParameterList objects.
3.7. Eigensolvers: Anasazi
Anasazi (Baker et al., 2009) is an extensible and interoperable framework for large-scale eigenvalue algorithms. This framework provides a generic interface to a collection of algorithms that are built upon abstract interfaces that facilitate code reuse and extensibility. Similar to Belos, Anasazi decouples the algorithms from the implementation of the underlying linear algebra objects using traits mechanisms. Concrete linear algebra adapters are provided for Tpetra and Thyra, while users can also implement their own interfaces to leverage any existing matrix and vector representations. Any libraries that understand Tpetra and Thyra matrices and vectors, like Belos and Ifpack2, may also be used in conjunction with Anasazi. The suite of eigensolvers provided by Anasazi includes locally optimal block preconditioned conjugate gradient (LOBPCG), block Davidson, Riemannian Trust-Region (RTR), and block Krylov-Schur methods. Recently, a family of trace minimization (TraceMin) methods and a generalized Davidson method have been added to the suite of eigensolvers in Anasazi.
3.8. Unified Solver Interface: Stratimikos
The Stratimikos package offers a cohesive interface to various linear solvers and preconditioners within Trilinos, including Amesos2, Belos, FROSch, Ifpack2, MueLu, and Teko. It requires that the matrix, right-hand side, and the solution vectors all support the Thyra interface. Thyra also provides wrappers for Tpetra linear algebra. Users can specify solver and preconditioner parameters via a Teuchos::ParameterList, which can be easily populated from an XML file. Moreover, the package allows for the easy addition of new solver and preconditioner factories via adapters, with examples available in the packages/stratimikos/adapters directory.
4. Nonlinear Solvers and Analysis Tools
The packages included in this section provide the top-level algorithms for computational simulations and design studies. These include nonlinear solvers, bifurcation tracking, stability analysis, parameter continuation, optimization, and uncertainty quantification. A common theme of this collection is the philosophy of “analysis beyond simulation”, which aims to automate many computational tasks that are often performed by application code users by trial-and-error or repeated simulation. Tasks that can be automated include performing parameter studies, sensitivity analysis, calibration, optimization, and the task of locating instabilities. Additional utilities for nonlinear analysis include abstraction layers and interfaces for application callbacks, as well as automatic differentiation tools that can provide the derivatives critical to analysis algorithms.
4.1. Nonlinear Solvers: NOX, LOCA
NOX (Nonlinear Object-oriented Solutions) provides robust and efficient algorithms for solving systems of nonlinear equations. NOX implements a number of Newton-based globalization techniques including line search (Pawlowski et al., 2006b), trust region (Pawlowski et al., 2006b, 2008), and homotopy algorithms (Coffey et al., 2003). Additionally, it provides lower- and higher-order models including Broyden, Anderson acceleration (Walker and Ni, 2011), and tensor methods (Bader, 2005). The algorithms have been designed for large-scale parallel inexact linear solvers using Krylov methods and support Jacobian-free Newton-Krylov variants. The library interacts with application codes through the Thyra model evaluator interface. NOX provides linear algebra abstractions for applications to use custom implementations with the algorithms. The library additionally contains abstractions for solvers, directions, and line searches that allow users to further customize the algorithms. NOX provides various stopping criteria, including absolute, relative, and weighted root mean square norms, as well as stagnation and NaN detection. Applications can build custom stopping criteria within a tree-based logical structure and provide additional criteria through the StatusTest abstraction. NOX also provides capabilities for continuation and stability analysis through the subpackage LOCA.
LOCA (Library of Continuation Algorithms) (Salinger et al., 2005) provides techniques for computing families of solutions to nonlinear equations as well as methods for investigating their stability when these nonlinear equations define equilibria of dynamical systems. It builds on the NOX nonlinear solver package to track solutions to sets of nonlinear equations as a function of one or more parameters (continuation). Given an interface to NOX that defines the nonlinear equations, users only need to provide a method for setting the parameter values being varied. LOCA provides several continuation methods, including pseudo-arclength continuation which allows tracking solution curves around turning points/folds. Furthermore, LOCA has hooks to call the Anasazi eigensolver package to estimate leading eigenvalues of the linearization at each point along the continuation curve for linear stability analysis, including various spectral transformations highlighting eigenvalues in different regimes (e.g., largest magnitude, largest real). Finally, LOCA implements equations augmenting the original nonlinear equations to locate and track bifurcation points where linear stability changes (e.g., turning point, pitchfork, and Hopf bifurcations) as a function of additional parameters. Examples of using LOCA for stability and bifurcation analysis include stability of a differentially heated cavity (Salinger et al., 2002), design of chemical vapor deposition reactors (Pawlowski et al., 2001), flow instabilities in counterflowing jets (Pawlowski et al., 2006a), and ignition in laminar diffusion flames (Shadid et al., 2006).
4.2. Numerical Optimization: ROL
Rapid Optimization Library (ROL) (rol, [n. d.]; Javeed et al., 2022) is the Trilinos package for numerical optimization. ROL brings an extensive collection of state-of-the-art optimization algorithms to virtually any application. Its programming interface supports any computational hardware, including heterogeneous many-core systems with digital and analog accelerators. ROL has been used with great success for optimal control, optimal design, inverse problems, image processing, and mesh optimization. Application areas include geophysics, climate science (Perego, 2022), structural dynamics (Aquino et al., 2019; Bunting et al., 2021), fluid dynamics (Antil et al., 2023b), electromagnetics, quantum computing, hypersonics, and geospatial imaging (Kouri and Shields, 2014).
The key design, performance, and algorithmic features of ROL can be summarized as follows:
  • Vector abstractions and matrix-free interface for universal applicability: Similar to other Trilinos packages that implement abstract numerical algorithms, ROL’s design is centered around an abstract linear algebra interface, through the ROL::Vector class, which enables the use of any ROL algorithm with any data type, such as the C++ std::vector, MPI-parallel data structures (e.g., Epetra, Tpetra, PETSc, HYPRE vectors), GPU-supporting data structures (e.g., Kokkos, ArrayFire (Yalamanchili et al., 2015)), etc. The abstract linear algebra interface further enables the definition of custom vector dot products, which are critical for performance in large-scale applications. Finally, all ROL algorithms are matrix-free and rely on user-defined applications of linear and nonlinear operators. ROL’s vector class design is inspired by the Hilbert Class Library (Gockenbach and Symes, 1996), the Rice Vector Library (Padula et al., 2009), and the RTOp framework (Bartlett et al., 2004), while ROL’s operator interfaces for simulation-based optimization, known as the SimOpt middleware, are inspired by (Heinkenschloss and Vicente, 1999).
  • Modern algorithms for unconstrained and constrained optimization: ROL features a large collection of well-established algorithms for smooth optimization as well as several novel algorithms developed by the ROL team. ROL categorizes optimization problems into four types based on the presence of certain constraints: Type U (unconstrained), Type B (bound constraints), Type E (equality constraints), and Type G (general constraints). Additionally, each problem type can include linear constraints, which ROL can reduce or eliminate algorithmically. The novel algorithms include augmented Lagrangian methods for infinite-dimensional problems with general constraints (Antil et al., 2023a) and matrix-free trust-region methods for convex-constrained optimization (Kouri, 2022).
  • Easy-to-use methods for stochastic and risk-aware optimization: ROL contains specialized interfaces and algorithms for stochastic optimization. Optimization problems with stochastic and uncertain problem data pose challenges in problem formulation and efficient numerical solution methods. ROL’s interface for stochastic optimization enables the use of numerous modern risk measures, probabilistic functions, and robust problem formulations. To approximate risk-averse problems (Shapiro et al., 2021), ROL includes sample-based approaches such as sample average approximation, stochastic approximation, and adaptive sparse-grid quadrature (Kouri et al., 2013, 2014). ROL also includes specialized algorithms such as progressive hedging (Rockafellar and Wets, 1991) and the primal-dual risk minimization algorithm (Kouri and Surowiec, 2022).
  • Fast and robust algorithms for nonsmooth optimization: ROL includes specialized algorithms to minimize the sum of a smooth function and a nonsmooth, convex function. This class of problems is ubiquitous in computational science. For example, nonsmooth sparsifying regularizers such as the $\ell^{1}$-norm are common in data science and machine learning. These algorithms can be used to solve convex constrained optimization problems. ROL’s methods for this class of problems are built on the proximity operator and include proximal gradient and Newton methods (Beck, 2017; Kanzow and Lechner, 2021; Ochs et al., 2014). As discussed in the subsequent bullet, ROL can also incorporate inexact values and gradients for the smooth objective function term using trust regions (Baraldi and Kouri, 2023). ROL also include an implementation of the trust bundle method (Schramm and Zowe, 1992) for solving general nonconvex, nonsmooth optimization problems.
  • Trust-region methods for inexact and adaptive computations: ROL controls inexact function evaluations and exploits adaptive computations and multi-fidelity models through its trust-region methods. Several of these methods have been pioneered by the ROL team. Their implementations typically require some knowledge of error in the computations (i.e., an error estimate). For a comprehensive overview of problem formulations, algorithms with error control, and their ROL implementations, see (Kouri and Ridzal, 2018).
  • PDE-OPT application development kit for PDE-constrained optimization: To build and optimize models based on partial differential equations (PDEs), ROL offers an application development kit called PDE-OPT. PDE-OPT is based on a modular design with three layers: local finite element computations, which encode the physical equations governing the optimization problem; global computations, which utilize Tpetra data structures; and an interface to ROL, enabling efficient optimization. To solve PDE-constrained optimization problems, users only need to implement the local finite element layer of PDE-OPT (i.e., the evaluations of weak forms on a mesh element); PDE-OPT automates all remaining steps.
4.3. Automatic Differentiation: Sacado
Sacado (Phipps and Gay, 2011; Phipps and Pawlowski, 2012; Phipps et al., 2008) provides forward and reverse-mode operator-overloading-based automatic differentiation (AD) tools within Trilinos. Sacado’s forward AD tools have been integrated with Kokkos and have demonstrated good performance on GPU architectures (Phipps et al., 2022). Sacado, along with its Kokkos integration, provides high-performance derivative capabilities to numerous Office of Science and NNSA (National Nuclear Security Administration) extreme scale applications, including: Albany for solid mechanics and land ice modeling (Salinger et al., 2016; Hoffman et al., 2018); Charon for semiconductor device modeling (Musson et al., 2020) and multiphase chemically reacting flows (Musson et al., 2009); Drekar for computational fluid dynamics (CFD) (Sondak et al., 2021; Shadid et al., 2016b), magnetohydrodynamics (Shadid et al., 2016a), and plasma physics (Crockatt et al., 2022; Miller et al., 2019); Xyce for electronic circuit simulation (Baker et al., 2012; Keiter et al., 2018); and SPARC for hypersonic fluid flows (Kieweg et al., [n. d.]).
4.4. Uncertainty Quantification: Stokhos
Stokhos (Phipps, 2015; Phipps and Salinger, 2016; Phipps et al., 2014) provides implementations of two intrusive uncertainty quantification strategies: the intrusive stochastic Galerkin uncertainty quantification method (Ghanem and Spanos, 1990, 2003) and the embedded ensemble propagation method (Phipps et al., 2017).
Stokhos’ implementation of the intrusive stochastic Galerkin uncertainty quantification method allows stochastic projections to be computed, such as polynomial chaos and generalized polynomial chaos expansions. The implementation includes interfaces for forming the linear and/or nonlinear systems that follow from the stochastic Galerkin projection, as well as linear solver methods that can exploit the block structure of these systems. GPU performance is targeted by using Kokkos and by commuting the layout of the Galerkin operator to be outer-spatial and inner-stochastic (Phipps et al., 2014). The implementation has been used in (Constantine et al., 2014) to efficiently propagate uncertainty in multiphysics systems by reducing the full system with a nonlinear elimination method.
The embedded ensemble propagation consists of propagating a subset of samples gathered into a so-called ensemble through the forward simulation at once. It builds on (Pawlowski et al., 2012a) for automating embedded analysis capabilities. Stokhos defines an ensemble type: a SIMD data type that is able to store the values of input, output, and state variables for every sample of an ensemble. This type can then be used within the Tpetra solver stack as a template argument for the scalar type. This approach allows to save computation time in four ways: the sample-independent data and computation can be reused for every sample of an ensemble, the memory access pattern is improved, the operations on the ensemble type can be vectorized efficiently, and the message passing costs are reduced by sending fewer but larger messages. However, the approach requires solvers and BLAS functions to be aware of the extra dimension associated to the ensemble; for example, a GMRES for ensemble types (Liegeois et al., 2020) needs to monitor the convergence of the individual samples in order to decide when to stop based on the union of the information.
4.5. Nonlinear Analysis Tools: Piro
Piro (Salinger and Phipps, 2010) provides a simple and unified interface to nonlinear analysis tools in Trilinos, exposing capabilities for computing sensitivities, performing parameter continuation and bifurcation analysis, and solving constrained optimization problems. In particular, Piro implements main driver classes for:
  • Linear/nonlinear solvers and sensitivity analysis for transient and nontransient problems. As an example, this capability can be used to compute the discrete solution of a partial differential equation depending on a parameter and the sensitivity of a quantity of interest with respect to this parameter. Sensitivities can be computed in a forward or adjoint fashion, the latter being preferable in the presence of high-dimensional parameters. This capability relies on NOX for the solution of nonlinear problems and Tempus for transient problems.
  • Constrained optimization problems with linear/nonlinear equality constraints: Piro provides tools for transient and nontransient (snapshot) optimizations, featuring gradient-based reduced-space and full-space methods. This capability is used by applications such as Albany to solve large-scale PDE-constrained optimization problems (Perego, 2022). Piro interfaces with ROL for algorithms to solve constrained optimization problems.
  • Parameter continuation and bifurcation analysis: this capability is provided through the package LOCA.
These driver classes share a similar interface based on the Thyra::ModelEvaluator and the Teuchos::ParameterList classes. Applications define the problems to be targeted (e.g., the equations to be solved, the parameters, the quantities of interests) by providing a concrete implementations of the Thyra::ModelEvaluator; see (Pawlowski et al., 2012a, b).
5. Discretization Tools
This section describes Trilinos packages that provide tools for spatial and temporal discretization of integro-differential equations. Most discretization efforts in Trilinos have been devoted to implementing tools for mesh-based discretizations of partial differential equations (PDEs) with a focus on high-order finite elements. A notable exception is the research package Compadre (Section 5.5), which provides tools for meshless approximation of linear operators that can be used for the discretization of differential equations and for data transfer. Trilinos’ discretization packages have been adopted by many applications addressing a wide range of physics problems, including solid mechanics, earth system modeling, semiconductor devices modeling, and electro-magnetics. These applications have taken different approaches in adopting Trilinos mesh-based discretization tools. The less intrusive approach is the adoption of Intrepid2 tools to perform local finite element assembly. In this case, the application has to manage the global assembly, possibly using the DoFManager provided by Panzer, and FECrs matrix and vector structures provided by Tpetra. A more intrusive approach is to additionally use the Phalanx package for managing dependencies of field evaluations in conjunction with the Thyra Model Evaluator and Sacado automatic differentiation, and possibly Tempus for time integration. This approach is particularly useful when developing solvers for complex multiphysics problems, because it allows easy re-use of computational kernels and automates the computation of Jacobians and sensitivities. The most intrusive approach is to build the application around the Panzer package, which provides all of the above, plus the handling of linear and nonlinear solvers and integrated constrained optimization capabilities. In the following, we neither include a description of the snapshotted packages Sierra ToolKit (STK) and Krino, which provide mesh and level-set tools, nor a description of the Shards package, which provides tools for mesh cell topology.
5.1. Local Assembly: Intrepid2
[…]
  • CellTools: This class provides geometric operations in the reference and physical frames. This includes the computation of tangents and normals to edges and faces in the physical frame, the computation of the Jacobian of the reference-to-physical frame map, and other geometric computations.
  • CubatureFactory: This class provides quadrature rules (called cubatures in Intrepid2) of various degrees of accuracy for approximating integrals over elements and their boundaries.
  • Basis: This is the base class that provides a common interface for functionalities related to finite element bases. Intrepid2 provides derived classes that implement this common interface for a variety of compatible finite element spaces. Each derived class implements the getValues() method that computes the values taken by the basis functions or their derivatives (e.g., the gradient for $H({\rm grad})$ functions, the curl for $H({\rm curl})$ functions) at a set of input points. The implementation of getValues() can be very different depending on the basis. Specific optimizations are available for tensor-product elements. Additionally, there is a BasisFamily class with a convenience method, getBasis(), which constructs a basis depending on a template argument specifying the type of basis (e.g., hierarchical or nodal), the cell topology, the function space on which it is defined, and its polynomial degree.
  • OrientationTools: This class provides methods to orient the basis functions based on the global orientation of edges and faces, determined by the global numbering of the cell vertices. This is achieved by building a linear operator (a permutation for tensor-product elements) that encodes the orientation of a particular cell and applying that operator to the reference basis functions.
  • ProjectionTools: This class provides methods for interpolation-based projections of a given function into a compatible finite element space or between compatible finite element spaces (Demkowicz et al., 2007). The provided projections commute with the corresponding differential operators if the quadrature rules can exactly integrate the functions being projected. As an example, projecting an $H({\rm grad})$ function into the $H({\rm grad})$ finite element space and then taking its gradient gives the same result as taking the gradient of the function first and then projecting the gradient into the $H({\rm curl})$ finite element space.
  • FunctionSpaceTools: This class provides transformations of fields from the reference to the physical frame and back, computation of measures on edges, faces, and cells, scalar/vector/tensor multiplications and contractions for computing integrals.
[…]
Intrepid2 makes use of Kokkos containers to enable memory layouts that are adapted to the computational platform. Intrepid2 also uses Kokkos for its core computational kernels, enabling threaded execution across a variety of architectures. The data types used by Intrepid2 are templated; it is therefore possible to propagate Sacado types through Intrepid2 to perform automatic differentiation. Current development of Intrepid2 focuses on providing efficient matrix-free discretizations to enhance efficiency on GPU architectures.
5.2. Local Field Evaluation: Phalanx
Phalanx is a local field evaluation library designed for equation assembly in PDE applications. The goal of Phalanx is to decompose a complex problem into a number of simpler problems with managed dependencies to support rapid development and extensibility of PDE codes (Notz et al., 2012; Pawlowski et al., 2012a, b). The data structures use Kokkos for performance portability. Through the use of template metaprogramming, Phalanx supports arbitrary user defined data types and evaluation types. This feature allows for simple integration of automatic differentiation tools via operator-overloaded scalar types from the Sacado package (Section 4.3). From a simple definition of equations, quantities such as Jacobians, Jacobian-vector products, Hessians, and parameter sensitivities can be evaluated to machine precision. These quantities can be used by other Trilinos packages for operations including Newton-based nonlinear solves, gradient-based optimization, constraint enforcement, and bifurcation analysis (Pawlowski et al., 2012a, b).
Phalanx uses a graph-based design to manage data dependencies. The runtime-defined directed acyclic graph (DAG) allows for rapid prototyping in a production environment where simple interfaces for analysts, flexible models/data structures, and integration of non-trivial third-party libraries are paramount. Phalanx is used in a number of large-scale parallel codes including Albany (Salinger et al., 2016), Charon (Musson et al., 2020), Drekar (Crockatt et al., 2022; Miller et al., 2019; Shadid et al., 2016a), and EMPIRE (Bettencourt et al., 2021).
In recent years, Phalanx has been extended to provide utilities for performance portability under automatic differentiation. For example, Phalanx provides tools for building and managing a Kokkos view-of-views on device without the use of unified virtual memory (UVM) and provides utilities for running virtual functions on device. In the future, these utilities may be separated into a separate package.
5.3. Time Integration: Tempus
Tempus (Latin, meaning time as in “tempus fugit” $\rightarrow$ “time flies”) is Trilinos’ time-integration package for advanced transient analysis. It includes various time integrators and embedded sensitivity analysis for next-generation code architectures. Tempus provides “out-of-the-box” time-integration capabilities, which allows users to quickly and easily incorporate time-integration capabilities in their applications and switch between various time integrators depending on the simulation needs. Additionally, Tempus provides “build-your-own” capabilities, which allows applications to incorporate various Tempus components to augment or replace an application’s transient capabilities. Other capabilities include embedded error analysis, sensitivity analysis, and transient optimization with ROL.
Tempus offers a general infrastructure for the time evolution of solutions through a variety of general integration schemes. Tempus provides time integrators for explicit and implicit methods and for first- and second-order ODEs. It can be used from small systems of equations (e.g., single ODEs for the time evolution of plasticity models and multiple ODEs for coupled chemical reactions) to large-scale transient simulations requiring exascale computing (e.g., flow fields around reentry vehicles and magneto-hydrodynamics).
Tempus has several components that can be used in concert or individually, depending on the needs of the application.
  • Integrators are the time-loop structure for time integration and provide several features, e.g., controlling the advancement of the solution, selecting the next time step size and handling the solution output.
  • Time steppers are individual methods that advance the solution from one step to the next. A variety of time steppers are available:
    • Classic one-step methods (e.g., forward Euler and trapezoidal methods)
    • Explicit Runge-Kutta (RK) methods (e.g., RK explicit 4 Stage)
    • Diagonally Implicit Runge-Kutta (DIRK) methods (e.g., general tableau DIRK and many specific DIRK/SDIRK methods)
    • Implicit-Explicit (IMEX) Runge-Kutta methods (e.g., IMEX RK SSP2, IMEX RK SSP3, and general tableau IMEX RK methods)
    • Multi-Step methods (i.e., BDF2)
    • Second-order ODE methods (e.g., Leapfrog, Newmark methods, and HHT-Alpha)
    • Steppers with subSteppers (e.g., operator-split and subcycling methods)
  • Solution history is used to maintain the solution during time step failure, for solution restart/output, for interpolation of the solution between time steps, and to provide the solution for transient adjoint sensitivities.
  • Time step control and strategies provide methods to select the time step size based on user input and/or temporal error control (e.g., bounding min/max time-step size, relative/absolute maximum error, and time step adjustments for output and checkpointing).
Additionally, Tempus has several mechanisms which allow users to insert application-specific algorithms into Tempus components (e.g., through observers and creating derived classes).
5.4. Finite Element Analysis: Panzer
The Panzer package provides global tools for finite element analysis. The package is targeted for large-scale parallel implicit PDEs using continuous and discontinuous compatible finite elements as well as control volume finite elements. Panzer enables the solution of nonlinear problems by interfacing with several Trilinos linear and nonlinear solvers. It uses the Intrepid2 package for arbitrary-order finite element bases. It computes derivatives and sensitivities with the Sacado automatic differentiation (AD) package. Panzer achieves performance portability through the Kokkos programming model via Tpetra data structures.
Panzer is designed for multiphysics systems. The assembly engine allows for different equation sets in each element block of the mesh and allows for mixed bases for degrees of freedom (DOF) within each element block (e.g., mixed $H(\rm{div})$ and $H(\rm{curl})$ system for electro-magnetics). The assembly tools can create a single fully coupled Jacobian matrix with all off-block dependencies (as a single Tpetra::CrsMatrix), or it can create a blocked system, grouping sets of DOFs into separate explicit matrices. This allows for physics-based block preconditioning strategies using the Teko package in Trilinos (Bonilla et al., 2023; Cyr et al., 2016) or MueLu’s fully coupled AMG hierarchies for multiphysics systems (Ohm et al., 2022). The assembly process relies on the Phalanx package for efficient assembly of the multiphysics systems. Panzer additionally can wrap the assembly in a Thyra::ModelEvaluator, providing direct interfaces to the linear and nonlinear analysis packages in Trilinos.
Panzer can provide low level utilities for application codes to build on, or can be used as a high-level application framework. Important capabilities include the following:
  • DOFManager: Panzer provides a stand-alone DOF manager class in the dof-mgr subpackage. Given a list of DOFs and their corresponding basis and element blocks, the DOF manager can provide the mapping from DOFs on the mesh entities to the entries in linear algebra objects such as residual vectors and Jacobian matrices. The DOF manager can return the objects required to build distributed Tpetra maps and graphs for both uniquely owned global indices and ghosted indices used during assembly. It provides the local indexing used during assembly as well. The DOF manager contains a mesh abstraction called a connection manager that provides information about mesh connectivity, global numbering of mesh topological entities, and element block groups. It is designed to support any underlying mesh database, allowing applications to use the DOF manager with any finite element application.
  • STK_Interface: Panzer contains a concrete implementation of the connection manager API for the STK mesh database package in Trilinos. The STK_Interface object wraps an STK mesh database. It provides a simple interface for accessing global indices of the mesh and can be used to read/write associated solution data to the database. It additionally supports SEACAS for writing mesh data to disk. The STK_Interface also includes support for periodic boundary conditions. This capability can match topological entities on periodic parallel distributed faces. Once matched, the DOFs are unified to enforce periodicity for the DOF manager. This capability is in the adapters-stk subpackage.
  • Linear Object Factory: Panzer provides a linear object factory and linear object container designed to support parallel distributed assembly. The returned containers hold either the linear algebra objects for a uniquely owned DOF map used for solving the linear system or a ghosted version of the linear objects that can be used for assembly. The containers support export and import operations between the unique and ghosted containers. These are used to simplify the assembly process and abstract the underlying Tpetra linear algebra objects. The linear object factory can also create DOF gather and scatter functors for the corresponding (and possibly blocked) matrices. The gather and scatter operations are specialized on the assembly types, including residuals, Jacobians, and tangents. Support for Hessians will be implemented as needed. The capability is found in the disc-fe subpackage.
  • Worksets: The Panzer assembly process provides workset containers to hold static data that can be reused in finite element computations. For example, when using a static mesh, these containers could hold the Jacobian, Jacobian inverse, and all finite element basis values at the quadrature points. The workset concept breaks the local elements on an MPI process into smaller blocks of uniform computations that can be dispatched to a CPU or GPU. The workset helps to control total memory use for temporaries in the Phalanx DAG. This capability is part of the disc-fe subpackage.
  • Assembly Kernels: Panzer provides a number of assembly kernels that are optimized for use with Sacado AD data types. When assembling on GPUs, the assembly kernels were found to be an order of magnitude faster if the AD evaluation was parallelized over the internal derivative dimension (Phipps et al., 2022). This required using kernels with a Kokkos::TeamPolicy for finite element assembly, where the lowest level of parallelism is used internally by the Sacado AD type. When building Panzer for GPUs, the DFad hierarchic parallelism flag in Sacado should be enabled for kernel performance. The kernels can be found in the disc-fe subpackage.
  • Panzer provides an example miniapp for implicit electro-magnetics that is used for benchmarking linear solver performance and for acceptance testing of new high-performance computing systems. It demonstrates an $H(\rm{curl})$-$H(\rm{div})$ formulation for the electric and magnetic fields. It can be found in the MiniEM subpackage.
The Panzer package is intended to provide both low- and high-level tools for implicit finite elements discretizations. The high-level tools aggregate many Trilinos discretization and solver packages. While the high-level tools can be used as a rapid prototyping environment, the front-end is fairly complex to set up, as opposed to true rapid prototyping frameworks such as deal.ii (Arndt et al., 2023), FEniCSx (Baratta et al., 2023), or Firedrake (Ham et al., 2023). Panzer is not user friendly in this regard, however, the examples and miniapps are a good starting point and can be quickly adapted to other physics applications. A number of performance portable applications are using Panzer tools at different levels of adoption; these include Albany (Salinger et al., 2016), Charon (Musson et al., 2020), Drekar (Crockatt et al., 2022; Miller et al., 2019; Shadid et al., 2016a), and EMPIRE (Bettencourt et al., 2021).
5.5. Approximation of Linear Operators: Compadre
The Compadre package provides tools for the approximation of linear operators (including point evaluation and derivatives), given the location of samples of a function over an unstructured cloud of points. The resulting stencils, when applied directly to samples of the function at these locations, provide an approximation of the linear operator acting on the function at the point(s) queried. This is useful for meshed and meshless data transfer applications (remap). Samples of the function at the specified locations can also be viewed as unknowns, in which case the solution returned by Compadre can be used as a stencil for meshless discretization of PDEs (Kuberry et al., 2024).
The package uses generalized moving least squares (GMLS) for approximating functionals. GMLS generalizes classic moving least squares (MLS) in the sense that selecting point evaluations for the sampling functionals and the target operator in GMLS provides a traditional moving least squares reconstruction. However, GMLS offers significantly greater flexibility than MLS by allowing the evaluation of the target function at points to be replaced with sampling functionals of a function (e.g., integrals of the function over local domains), and instead of approximating the target function, to approximate a target operator (e.g., gradient, curl, divergence) of a function. A detailed description of GMLS can be found in (Mirzaei et al., 2012; Wendland, 2004).
This increased flexibility enables more advanced and unconventional remapping approaches. For instance, it is possible to use an average vector normal integral over edges (Raviart-Thomas type representation) or a cell-average integral (finite-volume representation) as the sampling functionals or the target operator (Mahadevan et al., 2022). Compadre supports full space reconstruction in 1D, 2D, and 3D, and it also supports select sampling functionals and target operators on manifolds (Gross et al., 2020).
Compadre’s stencil generation involves independent problems to be solved in parallel at the team level with loops over the thread and vector level within each problem. This hierarchical parallelism is achieved with performance portability by using the Kokkos programming model and leveraging the batched QR with pivoting algorithm implemented in Kokkos Kernels.
Future plans include the implementation of other meshless methods like locally-supported radial basis functions.
6. Trilinos Framework
The Trilinos Framework product area provides supporting infrastructure for Trilinos users and developers. This infrastructure is largely not what Trilinos is well-known for but is essential for supporting all Trilinos capabilities.
6.1. Build and Test Infrastructure
The primary focus of the Trilinos Framework product relates to building, testing, and releases of Trilinos. This effort includes setting up and maintaining the general infrastructure for structured building of subsets of packages, automated testing, and maintaining and adding to the various testing configurations.
The Trilinos Framework infrastructure is built on top of the Tribal Build, Integration, and Test System (TriBITS (Bartlett, 2014)), which is built on top of the open-source tools CMake and CTest [^11] https://cmake.org. The TriBITS framework allows building arbitrary subgraphs of dependent (Trilinos) CMake packages in one or more individual aggregated CMake projects (in any arrangement desired). Each Trilinos/TriBITS package lists its direct (required and optional) dependent upstream packages, thus forming a package dependency graph. The TriBITS framework uses this package dependency graph to automatically determine what indirect dependent internal packages must be enabled and processed (and built) and what external packages must be found. TriBITS then orchestrates the processing of all of the required CMake code to find the needed external packages and configure and build (and optionally test and install) the selected set of internal packages. This allows a large number of (Trilinos) CMake packages to be configured, built, and tested in a flexible and efficient manner. In addition, TriBITS provides support for a number of advanced features that are not available in raw CMake/CTest including: eliminating a large amount of boiler-plate CMake code and avoiding common mistakes; enabling and testing all downstream packages given a set of enabled (i.e. modified) upstream packages; managing the enabling and disabling of tests based on various criteria; producing build and test results submitted to a CDash site on a package-by-package basis; producing reduced source tarballs for only a desired subset of enabled packages. As of version 14.4 of Trilinos, TriBITS and Trilinos have been updated to allow packages using raw CMake to be integrated with just a few well-defined integration requirements. The TriBITS framework has allowed Trilinos to scalably grow in the number of packages and the complexity without undue burden on individual Trilinos developers and users.
6.2. Documentation Infrastructure
The Trilinos Framework product is also responsible for maintaining the general infrastructure for the Trilinos website  [^12] https://trilinos.github.io, wiki  [^13] https://github.com/trilinos/Trilinos/wiki, and Doxygen documentation  [^14] https://trilinos.github.io/documentation.html. It owns some of the documented workflows on these sites, such as the process for reproducing testing failures, but the Trilinos package developers own all of the documentation associated with their associated packages.
6.3. Python Wrappers: PyTrilinos2
PyTrilinos2 is a set of automatically generated Python wrappers for selected Trilinos packages including Tpetra, Teuchos and Thyra, and for exposing solver capabilities from Amesos2, Belos, Ifpack2, and MueLu through Stratimikos. In the future, the list of wrapped packages will be enlarged to provide users with more features and to enable efficient prototyping of new algorithms for developers.
7. Trilinos Community
7.1. Contributing to Trilinos
Contributions to Trilinos can be offered through the standard GitHub pull request model. Proposed code changes are required to pass a set of tests as well as review and approval prior to be merged. Detailed instructions can be found in the contributing guidelines [^15] https://github.com/trilinos/Trilinos/blob/master/CONTRIBUTING.md in the source code repository.
7.2. Platforms for exchange among users or developers
The community of Trilinos users and developers operates several forums for exchange and discussion. Technical discussions about the source code and its development happen within the Trilinos GitHub repository [^16] https://github.com/trilinos/Trilinos. Several mailing lists (see https://trilinos.github.io/mail_lists.html) distribute relevant information and updates on Trilinos. The #trilinos channel within the Kokkos slack workspace [^17] See https://kokkos.org for details. provides a quick and accessible forum to ask questions.
For in-person exchange, the Trilinos User-Developer Group (TUG) Meeting takes place at Sandia National Laboratories in Albuquerque every year. At TUG, all Trilinos users and developers can gather to inform themselves about recent progress and advances, and discuss current challenges and upcoming topics relevant to the entire Trilinos community.
The European Trilinos User Group (EuroTUG) Meeting series [^18] https://eurotug.github.io offers a platform for Europe-based users and developers of the Trilinos project. EuroTUG offers Europe-based researchers and application engineers interested in the Trilinos project easy access to the Trilinos community, minimizing travel burdens by removing the need to travel to Albuquerque. It includes tutorials, user presentations showcasing Trilinos applications, and developer updates on news and ongoing work.
7.3. Embedding into other software initiatives
Trilinos is a founding member of the High Performance Software Foundation [^19] https://hpsf.io (HPSF). Established in 2024 under the Linux Foundation, the HPSF aims to build, promote, and advance a portable software stack for HPC by fostering collaboration among industry, academia, and government entities. As an initial technical project within the HPSF, Trilinos contributes its expertise in data structures for parallel computing, linear, nonlinear, and transient solvers, as well as optimization and uncertainty quantification in support of HPFS’ mission to enhance the HPC software ecosystem as a whole.
8. Concluding remarks
The continued existence and on-going development of the Trilinos project demonstrates its critical role in scientific and high-performance computing. This update underscores Trilinos’ dedication to maintaining relevance amidst the rapid development of HPC software frameworks and hardware architectures, and the demands of increasingly complicated multiscale multiphysics simulation code bases for tackling scientific and engineering problems. By adopting the Kokkos ecosystem, expanding package functionalities, and embracing a modular structure, Trilinos ensures both performance portability and adaptability to existing and emerging application domains.
The integration of data structures for distributed-memory paradigms, advanced linear and nonlinear solvers, discretization technology, and optimization tools highlights its versatility and impact across scientific and engineering disciplines. Its community-driven development model fosters innovation while ensuring robust support for users and contributors alike. Moving forward, the library’s commitment to collaboration, scalability, and innovation positions it as a cornerstone of next-generation computational frameworks.
As the computational landscape evolves, Trilinos remains poised to address emerging challenges, leveraging its rich feature set and strong community foundation to drive progress in scientific discovery and technological innovation.
Acknowledgments
Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA0003525.
SAND2025-02891O
This paper describes objective technical results and analysis. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government.
This work was supported in part by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Scientific Discovery through Advanced Computing (SciDAC) Program through the FASTMath Institute.
This work was supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Applied Mathematics program.
This work was supported by the Laboratory Directed Research and Development program at Sandia National Laboratories.
This work was supported in part by dtec.bw – Digitalization and Technology Research Center of the Bundeswehr [project hpc.bw]. dtec.bw is funded by the European Union – NextGenerationEU.